walnuts that cost $4.05 per kilogram were mixed w/ cashews that cost $7.25 per kilogram. How many kilograms of each were used to make a 50 kg. mixture that costs $6.25 per kilogram.(round to nearest 10th).
Answer by checkley77(12844)

7.25x+4.05(50-x)=6.25*50
7.25x+202.5-4.05x=312.50
3.2x=312.50-202.50
3.2x=110
x=110/3.2
x=34.38 kg. of cashews are used.
50-34.38=15.62 kg. of walnuts are used.
Proof:
7.25*34.38+4.05*15.62=312.50
249.26+63.26=312.50
-------------------------------------------------
Rounding to thew nearest pound would be 34 & 16.